A plate moves at 2 m/s with a shearing force of 3.5 N/m2. The distance between moving plate and fixed plate is 0.08mm. Determine viscosity of fluid between plates.



1
Expert's answer
2021-09-15T11:39:09-0400

The ratio of the shearing force to the area is


"\tau=\frac FA."

On the other hand, it is


"\tau=\mu\frac{u}{t}.\\\space\\
\tau=\mu\frac{u}{t},\\\space\\
\mu=\frac{\tau t}{u}=\frac{(3.5)(0.08·10^{-3})}{2}=1.4·10^{-4}\text{ Ns/m}^2."